OpenStudy (anonymous):

The cube root of 3 * the cube root of 9

OpenStudy (anonymous):

3

OpenStudy (anonymous):

could you explain?

OpenStudy (anonymous):

yes.... \[\sqrt[3]{3}*\sqrt[3]{9} = \sqrt[3]{27} = 3\]

OpenStudy (anonymous):

I thought you couldn't multiply radicals unless they have the same base?

OpenStudy (anonymous):

nope... that's not the case... :) as long as they are both cube roots... you can combine :) you can't add them though.. you can multiply

OpenStudy (anonymous):

ok thank you:) is it true that you can't multiply the fourth roots of 3 and -3? why is that if it follows what you said before?